OpenStudy (anonymous):

| 10-2i | I came up with sqrt 96 is that right?

OpenStudy (anonymous):

Nope. It will look like\[| 10-2i |=\sqrt{10^2+(-2)^2}=\sqrt{104}=2\sqrt{26} \approx 10.20\]

OpenStudy (anonymous):

ok after we get the 104 sqrt what do you do

OpenStudy (anonymous):

Remember, the absolute value is the distance from zero. With complex numbers, that is like the distance from the point that represents the number to the origin. Example: |a+bi| is represented on the complex plane by (a,b). Use the distance formula to find the distance between that point and (0,0).

OpenStudy (anonymous):

oh you just did the sqrt of 104. i see and you rounded it

OpenStudy (anonymous):

The simplification comes from the factoring of 104=2^3*13. Take the square root of 2^2, leave the rest inside the radical.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No sweat. Do math every day.